OpenStudy (spareb665):

In 2006, the population of a country was 40.7 million people. From 2006 to 2050, the countrys population is expected to decrease by 5.3%. Find the expected population of the country in 2050. Round to the nearest tenth of a million.

OpenStudy (anonymous):

decrease in population is 5.3% of 40.7 million = 2.2 million So, the population in 2050 is (40.7 million - 2.2 million) = 38.5 million

OpenStudy (spareb665):

hmm i got 0.053*40,700,000= 2157100 40,700,000-2157100=38542900

OpenStudy (anonymous):

5.3% of 40.7 = 5.3/100 x 40.7 = 2.1571 when rounded to nearest tenth, it becomes 2.2

OpenStudy (spareb665):

would you accept my answer as it is now?

OpenStudy (anonymous):

yes..you are right..but the answer should be rounded to nearest tenth of a million... 38542900 = 38.5492 million = 38.5 million when rounded

OpenStudy (anonymous):

Okay?

OpenStudy (spareb665):

ok Thanks alot!!!

OpenStudy (anonymous):

welcome : )